OpenStudy (anonymous):

Find the area of the region bounded by the graphs of y = −x^2 + 3x + 4 and y = 4.

OpenStudy (anonymous):

a. 2.7 b. 4.5 c. 28.5 d. 16.5

OpenStudy (anonymous):

@rational

OpenStudy (solomonzelman):

the area under the curve of the entire parabola is \(\large\color{slate}{\displaystyle\int\limits_{0}^{3}\left(-x^2+3x+4\right)~dx}\) and subtract the \(\large\color{slate}{\displaystyle\int\limits_{0}^{3}\left(4\right)~dx}\) from that

OpenStudy (solomonzelman):

Or, you can just shift the parabola 4 units down, and your area is \(\large\color{slate}{\displaystyle\int\limits_{0}^{3}\left(-x^2+3x\right)~dx}\)

OpenStudy (anonymous):

4.5 b

OpenStudy (anonymous):

@SolomonZelman can u help me on one more ?

OpenStudy (solomonzelman):

yes, 4.5 is correct

OpenStudy (solomonzelman):

I can try, I suck at math

OpenStudy (anonymous):

Find the area or the region bounded by the curves y = x^3 and y = 9x. a. 0 b. 10.13 c. 40.50 d. 20.25

OpenStudy (solomonzelman):

subtract the areas. first we need to solve for the second intersection point (besides x=0)

OpenStudy (solomonzelman):

(0,0) and (3,27) is the intersection (found it on the graph) where 9x >= x^3 on out interval of [0,3] so you can just subtract the areas

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{0}^{3}\left(9x\right)~dx-\int\limits_{0}^{3}\left(x^3\right)~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{0}^{3}\left(9x- x^3\right)~dx}\)

OpenStudy (solomonzelman):

on our (not out) interval ***

OpenStudy (anonymous):

20.25

OpenStudy (solomonzelman):

yes, correct